OpenStudy (anonymous):

area under curve y=0 x=y^4 y=(2-x)^(1/2)

OpenStudy (anonymous):

y=√(2-x) y² = 2-x 2-y² = x = y⁴ Let u = y² We find that y = ±1 So \(\displaystyle\int_{-1}^1 (2-y^2) - (y^4) \space dy\) Can you handle it now?

OpenStudy (anonymous):

why cant i integrate it using x with the bounds of 0 and 2

OpenStudy (anonymous):

You can do that, but doing it in terms of y is much easier. If you want to do it in term of x you have to break up area into two area and find integration separately.

OpenStudy (anonymous):

but it is bounded by y=0

OpenStudy (anonymous):

D'oh, I mean \[\int_0^1 (2- y^2) - (y^4)\space dy\] Either way, you will get the same answer.

OpenStudy (anonymous):

i got (16/5)(2^(1/4))-(4/3)(2^(1/2))

OpenStudy (anonymous):

None. You sure you set it correctly? \[\Large \int_0^1 \sqrt[4]{x} \space dx + \int_1^2 \sqrt{2-x}\space dx\]

OpenStudy (anonymous):

i didnt do that why do i have to split the integral?

OpenStudy (anonymous):